# Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations

### v1

@@ Abstract
We propose a probabilistic framework for developing computational models of biological neural systems. In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system which implements computations through its state evolution. To model this dynamical system, we employ a system of coupled stochastic differential equations with differentiable drift and diffusion functions and use variational inference to infer its states and parameters. This formulation enables seamless integration of existing mathematical models in the literature, neural networks, or a hybrid of both to learn and compare different models. We demonstrate this in our framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings. Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks show that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches while requiring an order of magnitude fewer parameters, providing uncertainty estimates, and offering a natural language for interpretation.
1 Introduction
@@
Unlike many scientific fields, neuroscience lacks a unified theoretical framework that bridges micro-scale mechanisms with macro-scale observations. That is, while the biophysical processes underlying individual neuron spiking have been well-understood for decades (Hodgkin & Huxley, 1952), the computational principles governing collective neural activity that give rise to cognition and behavior remain elusive.
@@
Dynamical systems theory provides a promising framework to address this gap (Durstewitz et al., 2023). Within this formalism, the brain is viewed as a complex dynamical system where neural activity patterns evolve through a high-dimensional state space according to well-defined governing equations, providing a mathematical language for understanding how distributed neural circuits implement computation and generate behavior (Van Gelder, 1998; Izhikevich, 2007; Deco et al., 2008; Breakspear, 2017; Favela, 2021). This perspective has led to diverse mathematical models, ranging from mean-field approximations of biophysical models that preserve mechanistic interpretability (Wilson & Cowan, 1972; Jirsa & Haken, 1996; Robinson et al., 1997), to phenomenological models inspired by statistical physics that capture emergent collective behavior (Amari, 1977; Hopfield, 1982; Sompolinsky et al., 1988). However, while these models have provided key insights into various neural phenomena (Goldman-Rakic, 1995; Wang, 2002; Fries, 2005), fitting them to single-trial neural data has proven challenging (Urai et al., 2022).
@@
In recent years, latent variable models have emerged as a powerful alternative. These models represent high-dimensional neural activity in terms of low-dimensional latent states evolving according to Markovian dynamics (Paninski et al., 2010; Hurwitz et al., 2021a). Neural computations are thus assumed to be implemented through dynamic motifs in a low-dimensional latent space (e.g., attractors, oscillations, bifurcations) (Vyas et al., 2020; Khona & Fiete, 2022; Durstewitz et al., 2023). The popularity of this approach stems from converging empirical evidence that neural population activity during simple tasks lives on a low-dimensional manifold (Gao & Ganguli, 2015). Coupled with ongoing developments in machine learning (Kingma & Welling, 2013; Fabius & Van Amersfoort, 2014; Rubanova et al., 2019), latent variable models are becoming a standard tool in systems and computational neuroscience (Chen et al., 2018; Glaser et al., 2020; Hurwitz et al., 2021b; Zhou & Wei, 2020; Kim et al., 2021).
@@
However, selecting an appropriate model for representing latent neural dynamics presents a significant challenge. The model must be expressive enough to capture complex non-linear dynamics, supporting flexible neural computations, while remaining amenable to interpretation and analysis to be useful for hypothesis generation and testing. Furthermore, model predictions should account for various sources of uncertainty typically present in such modeling endeavors, such as measurement noise, process noise, and model uncertainty. Finally, to utilize the inferred model for online applications as in brain computer interfaces (BCIs), there should be an efficient tractable method for sampling from the model in real-time.
@@
We posit that recent developments in scientific machine learning, where domain-specific models — often expressed as differential equations — are combined with neural networks and scalable optimization techniques provide an untapped potential to address these challenges (Raissi et al., 2019; Rackauckas et al., 2020; Lai et al., 2021; Karniadakis et al., 2021). Additionally, ongoing innovations in training (neural) stochastic differential equations (SDEs) (Tzen & Raginsky, 2019a; Li et al., 2020; Zeng et al., 2023; Course & Nair, 2024), provide an opportunity to develop scalable probabilistic models in neuroscience, a field wherein the systems in question are inherently stochastic (Laing & Lord, 2009; Rolls & Deco, 2010) and the existing measurement tools only provide a coarse proxy of the system under study (Stevenson & Kording, 2011; Urai et al., 2022).
Contributions
@@
We develop a general framework for modeling neural dynamics using latent SDEs with differentiable state- and input-dependent drift and diffusion functions. Within this framework, we introduce a specific generative model, combining coupled oscillators with neural networks, to capture neural population dynamics. We evaluate our approach across multiple experimental scenarios, demonstrating its effectiveness and efficiency in predicting stimulus-evoked neural and behavioral responses in both reaching and decision-making tasks. Through systematic comparison of latent ODE- and SDE-based models under varying noise levels and state dimensionality in simulated systems, we highlight the critical importance of modeling process noise; a factor often overlooked in current latent variable approaches. Finally, our results suggest that data-driven coupled oscillators could provide an expressive yet interpretable framework for understanding neural dynamics, while maintaining computational tractability.
2 Framework
Notations
@@ 2 Framework
Let denote a continuous path of latent neural states, which we observe through neural recordings (e.g., spike counts, fMRI time-series) denoted by , where each is observed at time . Additionally, we may observe behavioral responses (e.g., movement trajectories, choice responses) represented as , where each . In some settings, these observations are driven by external stimuli or task variables (e.g., visual stimuli, auditory cues) denoted by , where each . A single trial is thus defined as , where , and/or could be missing. Given a dataset consisting of trials, our objectives are to: (1) infer the underlying neural dynamics that generated the observations, represented by the posterior distribution , (2) learn a generative model capable of predicting stimulus-evoked responses, represented by the joint distribution , and (3) extract interpretable dynamical features that provide insights into neural computations.
2.1 Generative model
@@
This SDE induces a path measure on the space of continuous paths which implicitly define the distribution . Samples from this distribution are obtained by solving the SDE with different realizations of and :
@@
where and denote the observation models (e.g., Poisson model for spike counts, Gaussian model for continuous measurements), with and being differentiable functions mapping the latent state to the parameters of these distributions.
2.2 Inference and learning
@@
The true parameters of our generative model are unknown and need to be learned from the data . Moreover, the exact posterior distribution is intractable to compute. Variational inference provides an efficient and scalable approach to tackle this problem. However unlike standard latent variable models, the latent state in our model are realizations of a continuous-time stochastic process. Thus we need to define a parametric distribution over paths rather than random-valued vectors.
@@
Following (Tzen & Raginsky, 2019a, b; Li et al., 2020), we define the approximate posterior distribution via an augmented neural SDE. Specifically, here we define the approximate posterior path as realizations of the following SDE (we omit time indices for ease of notation):
where and is a continuous path encoding the observations, with being a modality-specific observation encoder. We define the distribution of the initial condition as , where and are neural networks, and . We further set to prevent over-parameterizing the initial conditions and rendering the dynamics redundant (Zabihi et al., 2019).
@@
Here is a neural network representing the drift of the augmented SDE. Note that while the drift term is different from that of the generative model, the diffusion term is shared. As illustrated in (Tzen & Raginsky, 2019a), this specification enables leveraging Girsanov’s theorem (Girsanov, 1960) to obtain a tractable KL divergence between the the generative (prior) and augmented (approximate posterior) distributions:
where with and denoting the measures induced by the generative and augmented SDE, respectively.
@@
Using this we can define an evidence lower bound (ELBO) on the conditional marginal likelihood of the observations:
@@
The ELBO consists of three main terms: (1) the expected log-likelihood of the observations under the approximate posterior, which includes both neural recordings and behavioral responses, (2) the KL divergence between the initial state distributions, and (3) the path-wise KL divergence between the approximate posterior and prior processes derived via Girsanov’s theorem (see Appendix B for detailed derivation of the loss).
The model parameters and variational parameters can be jointly optimized by maximizing this ELBO using stochastic gradient descent. Both the expectation over and the path-wise integral in the KL divergence term are computed using numerical integration schemes suitable for SDEs (see Appendix C for implementation details).
@@
This formulation enables us to simultaneously learn the parameters of the generative model while performing approximate posterior inference over the latent neural trajectories. The learned generative model can then be used for various online and offline downstream tasks including probabilistic prediction of stimulus-evoked responses and extraction of interpretable dynamical features.
3 Hybrid modeling
@@
So far we have described a general framework for modeling the latent dynamics of a neural system from empirical observations. This framework allows us to leverage almost arbitrary differentiable functions to model the drift and diffusion terms of an SDE. A natural choice would be to use a neural SDE to act as the generative model because of their universal approximation properties. However, this comes at the cost of efficiency and interpetability. An alternative option, is to leverage prior models proposed in the literature and use neural networks either to approximate their parameters or to augment their dynamics. Such hybrid models can improve sample-efficiency, generalization performance and, are easier to interpret in comparison to models that rely only on neural networks (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024). In this section, we detail a specific instantiation of the framework which combines a well-understood dynamical system and neural networks to obtain an expressive, interpretable and parameter-efficient generative model of neural population dynamics.
3.1 Stochastic nonlinear coupled oscillators
@@
A well-known class of models that are used to study several biological and physical systems are coupled oscillators. Coupled oscillators provide a rich yet simplified language for studying complex interactions in systems exhibiting emergent behavior such as synchronization, pattern formation and phase transitions (Winfree, 1980).
@@
We focus on a specific instantiation of coupled oscillators (Matthews et al., 1991) described by the following complex-valued ODE:
@@
Here we propose to use this system as the drift of our generative SDE in our framework. Specifically, we use the position of the oscillator as the latent state in the generative SDE. Additionally, we consider the coupling strength as a time-varying input dependent function of the external input .
Incorporating these updates and writing the model in terms of the real and imaginary parts of the oscillator , we obtain the following system of SDEs:
where , , are neural networks representing the input dependent connectivity and the diffusion terms for the real and imaginary states respectively. The trainable parameters of the generative SDE are thus the parameters of the neural networks as well as and . We refer to this dynamics model as coupled oscillator SDE (CO-SDE).
3.2 Modeling multiple interacting populations
Often, we have access to neural activity recorded from neurons across multiple brain regions, simultaneously. Given access to recordings from brain regions, we can extend the generative model by factorizing the latent state where the number of oscillators can vary depending on the population size and complexity. The observed activity of a population is then
@@ 3.2 Modeling multiple interacting populations
4 Experiments
4.1 Simulated data
We first sought to evaluate the efficacy of our framework on modeling a partially observed neural system with different levels of process noise. To this end, we generated spiking data from the following system:
where describes the population average firing rate of three neural populations driven by a sinusoidal input with randomly sampled frequency and phase. and describe the recurrent connectivity and input interaction matrices, respectively. The time scale of the network dynamics is given by and describes the magnitude of process noise in the system. The system output is given by representing the observed spike counts of 50 neurons generated by sampling from an inhomogeneous Poisson process with rates computed by affine transformation of the population average activity via the output matrix . We fit our latent SDE model on the spiking data by maximizing the Poisson-log likelihood of the observations given predicted rate and evaluated performance using bits per spike (bps), a metric that quantifies how well the predicted firing rates explain the observed spike patterns (Pei et al., 2021).
For our experiments, we varied the noise level to test the robustness of the model. We used a latent neural SDE with varying sizes of latent dimension and compared against two baselines: a latent ODE (Rubanova et al., 2019) and LFADS (Pandarinath et al., 2018).
As demonstrated in Figure 2a, the latent SDE model can better explain the data variance when using fewer latent dimensions compared to a latent ODE model. Increasing the latent dimensions beyond the true state dimensions does not significantly improve the model fit unlike the ODE-based model which stagnates at a lower absolute value and higher latent dimension. Additionally, as shown in Figure 2b, the latent SDE model is more robust to process noise compared to latent ODE and LFADS models. These results suggest that latent SDEs can be more reliable to model real-world data when the true state dimension is unknown and process noise can emerge from unobserved interactions. In this setting, a probabilistic initial state with deterministic dynamics, as in the case of latent ODEs and LFADS, may not be sufficient.

### v2

@@ Abstract
Understanding how neural circuits implement computation remains a central challenge in neuroscience, requiring models that can unite mechanistic interpretability with the complexity of real neural data. Current approaches face a fundamental trade-off: mechanistic biophysical and phenomenological models are difficult to fit to single-trial data, while flexible machine learning models lack interpretability and theoretical grounding. We propose a probabilistic framework that addresses this challenge by combining domain-specific mathematical models with neural networks through stochastic differential equations (SDEs). In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system that implements computations through its state evolution. We employ coupled SDEs with differentiable drift and diffusion functions and use variational inference to infer system states and parameters. This formulation enables seamless integration of existing mathematical models from the literature, neural networks, or hybrid combinations to learn and compare different modeling approaches. We demonstrate this framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings. Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks shows that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches, while requiring an order of magnitude fewer parameters, providing principled uncertainty estimates, and offering interpretable mathematical descriptions of neural dynamics.
Introduction
@@
Unlike many scientific fields, neuroscience lacks a unified theoretical framework that bridges micro-scale mechanisms with macro-scale observations. That is, while the biophysical processes underlying individual neuron spiking have been well-understood for decades [33], the computational principles governing collective neural activity that give rise to cognition and behavior remain elusive.
@@
Dynamical systems theory provides a promising framework to address this gap [60, 18]. Within this formalism, the brain is viewed as a complex dynamical system, where neural activity patterns evolve through a high-dimensional state space according to well-defined governing equations, providing a mathematical language for understanding how distributed neural circuits implement computation and generate behavior [78, 38, 15, 9, 23]. This perspective has led to diverse mathematical models, ranging from mean-field approximations of biophysical models that preserve mechanistic interpretability [83, 39, 63] to phenomenological models inspired by statistical physics that capture emergent collective behavior [3, 34, 70]. However, while these models have provided key insights into various neural phenomena [31, 82, 25], fitting them to single-trial neural data has proven challenging [76].
@@
In recent years, latent variable models have emerged as a powerful alternative. These models represent high-dimensional neural activity in terms of low-dimensional latent states evolving according to Markovian dynamics [58, 36]. Neural computations are thus assumed to be implemented through dynamic motifs in a low-dimensional latent space (e.g., attractors, oscillations, bifurcations) [81, 41, 18]. The popularity of this approach stems from converging empirical evidence that neural population activity during simple tasks lives on a low-dimensional manifold [28]. Coupled with ongoing developments in machine learning [45, 22, 65], latent variable models are becoming a standard tool in systems and computational neuroscience [11, 30, 37, 89, 43].
@@
However, selecting an appropriate model for representing latent neural dynamics presents a significant challenge. The model must be expressive enough to capture complex non-linear dynamics, supporting flexible neural computations, while remaining amenable to interpretation and analysis to be useful for hypothesis generation and testing. Furthermore, model predictions should account for various sources of uncertainty typically present in such modeling endeavors, such as measurement noise, process noise, and model uncertainty. Finally, to utilize the inferred model for online applications, as in brain-computer interfaces, there should be an efficient tractable method for sampling from the model in real time.
@@
We posit that recent developments in scientific machine learning – where domain-specific models, typically expressed using differential equations, are combined with neural networks and scalable optimization techniques – provide an untapped potential to address these challenges [62, 61, 47, 40]. Additionally, ongoing innovations in training (neural) stochastic differential equations (SDEs) [74, 49, 87, 14], provide an opportunity to develop scalable probabilistic models in neuroscience, a field wherein the systems in question are inherently stochastic [48, 64] and the existing measurement tools only provide a coarse proxy of the system under study [72, 76].
@@
To make these advances in scientific machine learning available to the neuroscience community, we developed a general framework for modeling neural dynamics using latent SDEs with differentiable state- and input-dependent drift and diffusion functions. Within this framework, we introduce a specific generative model, combining coupled oscillators with neural networks, to capture neural population dynamics. We evaluate our approach across multiple experimental scenarios, demonstrating its effectiveness and efficiency in predicting stimulus-evoked neural and behavioral responses in both reaching and decision-making tasks. Through systematic comparison of latent ODE- and SDE-based models under varying noise levels and state dimensionality in simulated systems, we highlight the critical importance of modeling process noise; a factor often overlooked in current latent variable approaches. Finally, our results suggest that data-driven coupled oscillators could provide an expressive yet interpretable framework for understanding neural dynamics, while maintaining computational tractability.
Related work
Latent variable models in neuroscience
Our work builds on previous latent variable models which have been proposed to infer low-dimensional latent dynamics from single-trial data [68, 86, 50, 85, 57, 16, 37, 69, 43, 44]. Most related to our work, is Gaussian process switching linear dynamical system [2] which leverage SDEs to represent latent neural dynamics extending previous work proposed by Duncker et. al [16]. In their work, the authors developed a novel Gaussian process kernel function that defines a smooth, locally linear prior on dynamics, yielding easily interpretable recurrent linear switching dynamical systems while providing posterior uncertainty estimates. Compared to Gaussian process-based methods, our approach provides flexibility by allowing the modeler to choose any dynamical system as the prior, enabling systematic model comparison across a wide range of existing neural models. Additionally, our approach is not limited to a small number of latent state parameters as is often the case with GP-based approaches, enabling greater flexibility in model capacity. Finally, unlike [2], the SDEs employed in our approach are not restricted to additive process noise, but can model multiplicative (state- and input-dependent) process noise, enabling greater flexibility and model expressivity. Other relevant recent work extends low-rank RNNs to a stochastic setting with additive process noise [56]. Here, the authors use a specific form of the RNN which enables finding fixed-points effectively after training.
Scalable inference of latent SDEs
Inferring the state and learning the parameters of a latent SDE from (high-dimensional) noisy observations is a challenging problem which appears throughout science and engineering. With the advent of neural differential equations [11, 42], there has been a growing interest in developing scalable methods for fitting latent neural SDEs to data. Seminal work by [74, 75] established the theoretical foundation for training generative neural SDEs via variational inference through the lens of stochastic control. Subsequent research has focused on addressing key computational challenges, including improved memory efficiency and extension to the time domain [49], enhanced numerical stability [87, 55], and computational efficiency [14]. Our work builds upon the formulation by [49], extending it to controlled and multi-modal settings in neuroscience. This specific formulation enables preserving the expressivity of the generative SDE through state- and input-dependent diffusion on arbitrary spaces.
Modeling neural dynamics with coupled oscillators
Coupled oscillator models have a rich history in theoretical neuroscience, dating back to seminal work by [46] and [83]. These models have proven particularly effective in capturing rhythmic neural activity and investigating synchronization phenomena in neural circuits [8, 21]. Of particular relevance to our work are modern approaches that combine oscillator models with machine learning techniques [1, 2, 4] to learn dynamical systems that can reproduce complex neural trajectories while maintaining interpretable structure. Our approach builds upon these foundations by introducing a flexible SDE framework that can naturally capture oscillatory dynamics while accounting for process noise and external inputs. This bridges the gap between mechanistic oscillator models and data-driven approaches, allowing for both accurate trajectory prediction and meaningful scientific interpretation.
Methods
Notation
@@ Methods
Let denote a continuous path of latent neural states, which we observe through neural recordings (e.g., spike counts, fMRI time-series) denoted by with is observed at time . Additionally, we may observe behavioral responses (e.g., movement trajectories, choice responses) represented as with . In some settings, these observations are driven by external stimuli or task variables (e.g., visual stimuli, auditory cues) denoted by with . A single trial is thus defined as , where , and/or could be (partially) missing. Given a dataset consisting of trials, our objectives are to: (1) infer the underlying neural dynamics that generated the observations, represented by the posterior distribution , (2) learn a generative model capable of predicting stimulus-evoked responses, represented by the joint distribution , and (3) extract interpretable dynamical features that provide insights into neural computations.
Generative model
@@
This SDE induces a path measure on the space of continuous paths which implicitly define the distribution . Samples from this distribution for are obtained by solving the SDE with different realizations of and :
@@
where and denote the observation models (e.g., Poisson model for discrete measurements, Gaussian model for continuous measurements) for neural and behavioral data, respectively, with and being differentiable functions mapping the latent state to the parameters of these distributions.
Model estimation
@@
The true parameters of our generative model are unknown and need to be learned from data . Moreover, the exact posterior distribution is intractable to compute. Variational inference provides an efficient and scalable approach to tackle this problem. Unlike standard latent variable models, the latent state in our model are realizations of a continuous-time stochastic process. Thus we need to define a parametric distribution over paths rather than random-valued vectors. Figure 1 provides an overview of our framework.
@@
Following [74, 75, 49], we define the approximate posterior distribution via an augmented (neural) SDE. Specifically, we define the approximate posterior path as realizations of the following SDE (we omit time indices for ease of notation):
where and is a continuous path encoding the observations which we refer to a context, with being a modality-specific observation encoder. We define the distribution of the initial condition as , where and are neural networks, and are discrete samples of the continuous-time context . Here, is a neural network representing the drift term of the augmented SDE. Note that while the drift term is different from that of the generative model, the diffusion term is shared with the generative SDE.
@@
As described in [74], this specification enables leveraging Girsanov’s theorem [29] to obtain a tractable KL divergence between the generative (prior) and augmented (approximate posterior) distributions:
@@
where with and denoting the measures induced by the generative and augmented SDE, respectively. Using this property, we can define the following evidence lower bound (ELBO) on the conditional marginal likelihood of the observations:
@@
The ELBO consists of three main terms: (1) the expected log-likelihood of the observations under the approximate posterior, which includes both neural recordings and behavioral responses, (2) the KL divergence between the initial state distributions, and (3) the path-wise KL divergence between the approximate posterior and prior processes derived via Girsanov’s theorem (see Supplementary Information for a detailed derivation of the loss). The model parameters and variational parameters can be jointly optimized by maximizing this ELBO using stochastic gradient descent. Both the expectation over and the path-wise integral in the KL divergence term are computed using numerical integration schemes suitable for SDEs (see Supplementary Information for implementation details).
@@
This formulation enables us to simultaneously learn the parameters of the generative model while performing approximate posterior inference over the latent neural trajectories. The learned generative model can then be used for various online and offline downstream tasks, including probabilistic prediction of stimulus-evoked responses and extraction of interpretable dynamical features.
Biophysics-inspired latent SDE models
@@
So far, we have described a general framework for modeling the latent dynamics of a neural system from empirical observations. This framework allows us to leverage almost arbitrary differentiable functions to model the drift and diffusion terms of an SDE. A natural choice would be to use a neural SDE to act as the generative model because of their universal approximation properties. However, this comes at the cost of efficiency and interpretability.
An alternative option, is to leverage prior models proposed in the literature and use neural networks either to approximate their parameters or to augment their dynamics. Such hybrid models can improve sample-efficiency and generalization performance, and are easier to interpret in comparison to models that rely only on neural networks [61, 20].
In this section, we detail a specific instantiation of the framework which combines a well-understood dynamical system and neural networks to obtain an expressive, interpretable and parameter-efficient generative model of neural population dynamics. Furthermore, we describe how such models can be linked to neurobehavioral data via the specification of appropriate forward models.
Stochastic nonlinear coupled oscillators
@@
A well-known class of models that are used to study several biological and physical systems are coupled oscillators. Coupled oscillators provide a rich yet simplified language for studying complex interactions in systems exhibiting emergent behavior such as synchronization, pattern formation and phase transitions [84].
@@
We focus on a specific instantiation of coupled oscillators [53] described by the following complex-valued ordinary differential equation (ODE):
@@
Here we propose to use this system as the drift of our generative SDE in our framework. Specifically, we use the position of the oscillator as the latent state in the generative SDE. Additionally, we consider the coupling strength as a time-varying input-dependent function of the external input . Incorporating these updates and writing the model in terms of the real and imaginary parts of the oscillator , we obtain the following system of SDEs:
where represents the state of coupled oscillators and , , are neural networks representing the input dependent connectivity and the diffusion terms for the real and imaginary states respectively. The trainable parameters of the generative SDE are thus the parameters of the neural networks as well as and . We refer to this dynamics model as a coupled-oscillator SDE (CO-SDE).
Modeling neurobehavioral observations
Often, we have access to observational data, which can come in the form of neural activity recorded from multiple brain regions or in the form of behavioral measurements. To inform the latent SDE model, we need to condition the generative model on such observational data.
Given access to neural activity recorded from brain regions, we can extend the generative model by factorizing the latent state , where the number of oscillators can vary depending on the population size and complexity. Assuming that the neural data comes in the form of population firing rates, we model the observed activity of a population as
@@
For behavioral observations, in case of continuous-valued data (e.g. arm position, pupil diameter), we assume a Gaussian measurement model:
where and are neural networks that map the latent state to the mean and variance parameters of the observation distribution, respectively.
Finally, for discrete-valued behavioral data (e.g., choice response), we employ a categorical observation model:
where is a neural network followed by a softmax activation that maps the latent state to a probability vector over discrete categories.
Note that neural and behavioral data can be easily combined by using the described forward models in parallel. This provides a natural solution to the data fusion problem and provides a principled way to integrate multi-modal data in our analyses.
Results
Validation on simulated data
We first sought to evaluate the efficacy of our framework for modeling a partially observed neural system with different levels of process noise. To this end, we generated spiking data from the following system:
where describes the population average firing rate of three neural populations driven by a sinusoidal input with randomly sampled frequency and phase. The system matrix and input matrix describe the recurrent connectivity and input dependence, respectively. The time scale of the network dynamics is given by and describes the magnitude of process noise in the system. The system output is given by representing the observed spike counts of 50 neurons generated by sampling from an inhomogeneous Poisson process with rates computed by affine transformation of the population average activity via the output matrix .
We fit our latent SDE model on the spiking data by maximizing the Poisson-log likelihood of the observations given predicted rate and evaluated performance using bits per spike (bps), a metric that quantifies how well the predicted firing rates explain the observed spike patterns [59]. For our experiments, we varied the noise level to test the robustness of the model. We used a latent neural SDE with varying sizes of latent dimension and compared against two baselines: a latent ODE [65] and LFADS [57] as these have gained popularity in the field.
As demonstrated in Fig. 2a, the latent SDE model can better explain the data variance when using fewer latent dimensions compared to a latent ODE model. Increasing the latent dimensions beyond the true state dimensions does not significantly improve the model fit unlike the ODE-based model which stagnates at a lower absolute value and higher latent dimension. Additionally, as shown in Fig. 2b, the latent SDE model is more robust to process noise compared to latent ODE and LFADS models. These results suggest that latent SDEs can be more reliable when modeling real-world data when the true state dimension is unknown and process noise can emerge from unobserved interactions. In this setting, a probabilistic initial state with deterministic dynamics, as in the case of latent ODEs and LFADS, may not be sufficient. Figures 2c and 2d show the model predictions for the spike counts and the underlying firing rates, respectively. Figure 2e demonstrates the phase portrait of the inferred dynamical system compared to the ground truth simulated dynamical system.
Validation on empirical data